OpenStudy (anonymous):

prove that cosh(2x) = 2sinh^2(x) + 1 *Recall that sinh(x) = (e^x - e^-x)/2 and cosh(x) = (e^x + e^-x)/2... Also, cosh^2(x) - sinh^2(x)=1 Can anyone help me?

OpenStudy (anonymous):

substitute and prove LHS = RHS

OpenStudy (anonymous):

\[2(\frac{e^x-e^{-x}}{2})^2+1=(\frac{e^{2x}-2+e^{-2x}}{2})+1=\] make it one fraction by adding 1

OpenStudy (anonymous):

Aani, I know that I am suppose to prove that the left equals the right...that is what a proof means...anything else you can contribute? Thanks

OpenStudy (anonymous):

Hi Jonask...let me take a second and read over your post...thanks

OpenStudy (anonymous):

\[\frac{e^{2x}+e^{-2x}}{2}=\cosh 2x\]

OpenStudy (anonymous):

@SinginDaCalc2Blues so first we put wat sinh is in the right hand side equation

OpenStudy (anonymous):

HI again Jonask...I am trying to read your equations but I think the computer changed their format and I simply can't understand them...do you see what I see?

OpenStudy (anonymous):

okay i'll send a picture of them

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

I have tried solving down both sides many times...it looks like the right side might be better to manipulate but I'm not certain...what do you think?

OpenStudy (anonymous):

Hey Jonask...I got it!!! Thank you! I was able to figure out the jibberish equations! Awesome! Thank u!

OpenStudy (anonymous):

here you go took some time

OpenStudy (anonymous):

yeah i thought i had it but i still had a 2 in front...one second and i will look at your work...

OpenStudy (anonymous):

the two went out because the denominator was 2^2=4 so canceling the two on top

OpenStudy (anonymous):

Amazing work!...That's what I was missing, the darn 4 in the bottom! haha Thank yiou soooo much for youe effort!!! I would give yoiu 2 gold medals if I could!

OpenStudy (anonymous):

now some other easy way lets consider the right hand side again\[2\sinh^2x+1=2\sinh^2x+\cosh^2x-\sinh^2x=\sinh^2x+\cosh^2x=\cosh2x\]

OpenStudy (anonymous):

@SinginDaCalc2Blues

OpenStudy (anonymous):

going through your 2nd solution now...

OpenStudy (anonymous):

I dont follow this last part, if I'm reading it correctly... \sinh^2x+\cosh^2x=\cosh2x\] Is that saying: (sinh^2x + cosh^2x) = cosh^2x ??? If I'm reading that correctly, I don't understand

OpenStudy (anonymous):

I gave you a medal and became your fan! If you happen to write anything else, I will come back here and check it out...thanks again!

OpenStudy (anonymous):

Hey Jonask, if you feel like helping me one more time, I posted a new question...if you have the time

OpenStudy (anonymous):

\[\sinh^2x+\cosh^2x=(\frac{e^x-e^{-x}}{2})^2+(\frac{e^x+e^{-x}}{2})^2=\cosh 2x \] i notice that for the last part yuo have to show this 1st